OpenStudy (anonymous):

find the volume of a sphere with a diameter of 22 inches.use 3.14 to approximate n. round answere to nearest hundreth.

OpenStudy (anonymous):

Do you know the equation for the volume of a sphere?

OpenStudy (callisto):

Radius r = diameter /2 = 22/2 =11 Put \(\pi = 3.14\)and r= 11 into the formula \(V=\frac{4}{3} \pi r^3\) to find the volume.

OpenStudy (anonymous):

i got 5572.453

OpenStudy (callisto):

Yes, but you need to '' round answer to nearest hunredth.''

OpenStudy (anonymous):

so it will = 60123.41

OpenStudy (callisto):

Why?

OpenStudy (anonymous):

i thought my answer was rounded .

OpenStudy (callisto):

Hm.. but that's not the answer. The answer you've got was 5572.453. That sounds better. The only problem is that you haven't rounded this answer.

OpenStudy (anonymous):

when i rounded i got 5572

OpenStudy (anonymous):

Rounding to the nearest hundredth means 5572.45. You rounded to the nearest whole number.

OpenStudy (anonymous):

ok thanks.